OpenStudy (anonymous):

find the interecepts for the graph of the equation given. 2x-4y=12 The x-intercept is ? The y-intercept is ?

jimthompson5910 (jim_thompson5910):

To find the x-intercept, plug in y = 0 and solve for x 2x-4y=12 2x-4(0)=12 2x - 0 = 12 2x = 12 x = 12/2 x = 6 I'll let you find the y-intercept. You do so by plugging in x = 0 and solving for y.

OpenStudy (anonymous):

-3?

jimthompson5910 (jim_thompson5910):

good, that's the y-intercept